the length of a rectangle is 3 feet less than twice the width. If the area is 77 sq feet, then solve a quadratic equation to find the length and width.
Answer by fractalier(6550)

From the info in the problem,
L = 2W - 3
We know the formula for the area is
A = LW and it equals 77, so now substitute in for L and solve:
A = LW = 77
(2W - 3)W = 77
2W^2 - 3W - 77 = 0
(2W + 11)(w - 7) = 0
W = -11/2 or W = 7
Length can't be negative, so W = 7.
L = 2W - 3 = 2(7) - 3 = 11